OpenStudy (anonymous):

What value of b makes this equation true? 5/9 . 9/b = 1 A. –5 B. 0 C. 5 D. 1

OpenStudy (anonymous):

divide both sides by 5/9 when you divide by a fraction you multiply by its reciprocal so 9/b=1*(9/5) 9/b=9/5 should be obvious now

OpenStudy (anonymous):

thnx :)